# How Data Mixing Shapes In-Context Learning: Asymptotic Equivalence for Transformers with MLPs

### v1

@@ Appendix B Asymptotic distribution of ${\bm{F}}\text{vec}({\bm{H}}_{\bm{Z}})$
$\delta_{x}:=\frac{\left|\|{\bm{x}}_{\ell+1}\|^{2}-\text{Tr}(\text{Cov}({\bm{x}}_{\ell+1}))\right|}{\text{Tr}(\text{Cov}({\bm{x}}_{\ell+1}))},\quad\text{and}\quad\delta_{v}:=\frac{\left|\|{\bm{b}}\|^{2}-\text{Tr}(\text{Cov}({\bm{b}}))\right|}{\text{Tr}(\text{Cov}({\bm{b}}))}.$ (S16)
@@
$\frac{\left|\|\text{vec}({\bm{H}}_{\bm{Z}})\|^{2}-\text{Tr}(\text{Cov}(\text{vec}({\bm{H}}_{\bm{Z}})))\right|}{\text{Tr}(\text{Cov}(\text{vec}({\bm{H}}_{\bm{Z}})))}$ $=\frac{\left|\|{\bm{x}}_{\ell+1}\|^{2}\|{\bm{b}}\|^{2}-\text{Tr}(\text{Cov}({\bm{x}}_{\ell+1}))\text{Tr}(\text{Cov}({\bm{b}}))\right|}{\text{Tr}(\text{Cov}({\bm{x}}_{\ell+1}))\text{Tr}(\text{Cov}({\bm{b}}))}$ (S17)
@@
$t=\text{Tr}(\text{Cov}(\text{vec}({\bm{H}}_{{\bm{Z}}})\mid s=i))=\text{Tr}(\text{Cov}(\text{vec}({\bm{H}}_{{\bm{Z}}})\mid s=j))$
for any data source indices $i,j$. Furthermore, Corollary B.2 showed that $\|\text{vec}({\bm{H}}_{\bm{Z}})\|^{2}/t$ concentrates around 1 for all $s$ and $\bm{\xi}|s$. This leads us to the following unconditioned concentration result.
@@
$\left|\frac{\|\text{vec}({\bm{H}}_{{\bm{Z}}})\|^{2}}{t}-1\right|\leq\max_{s,\bm{\xi}|s}\left|\frac{\|\text{vec}({\bm{H}}_{{\bm{Z}}})\mid(s,\bm{\xi}|s)\|^{2}}{t}-1\right|=o(1),$ (S20)
@@
We apply a union bound over the (countable) set of data sources and integrate over the distribution of $\bm{\xi}|s$. Since the conditional deviation is uniformly bounded by $o(1)$, the same bound extends to the unconditioned case. ∎
@@
This completes our analysis of the concentration of $\|\text{vec}({\bm{H}}_{\bm{Z}})\|^{2}/t$ around 1, and thereby concludes the proof of Lemma 4.9.
@@ Appendix C Decomposition of the gradient matrix
where $\alpha:=\mathbb{E}_{z\sim\mathcal{N}(0,1)}[\sigma^{\prime}(z)]$ is the average slope, and $\sigma^{\prime}_{\perp}(z):=\sigma^{\prime}(z)-\alpha$ captures the zero-mean deviation. The decomposition is justified by the fact that the random variable entering $\sigma$—namely, ${\bm{f}}_{i}^{T}\tilde{{\bm{h}}}_{j}$—converges in distribution to $\mathcal{N}(0,1)$ almost surely for all $i,j$, by Lemma 4.9.
@@
This decomposition separates the gradient into two components: A spike term ${\bm{u}}{\bm{v}}^{T}$, which is expected to dominate, and a residual term $\bm{\Delta}$, which we aim to show is negligible.
Our next goal is to establish that $\|{\bm{u}}{\bm{v}}^{T}\|\gg\|\bm{\Delta}\|$, thereby confirming that the spike term governs the spectral structure of $\bm{G}$}.
@@
$\|\tilde{{\bm{H}}}\|\,/\,\|\mathrm{Cov}(\text{vec}({\bm{H}}_{{\bm{Z}}}))\|^{1/2}={\tilde{{\mathcal{O}}}}(d).$ (S23)
@@
$\left|\|\text{vec}({\bm{H}}_{{\bm{Z}}})\|^{2}-\operatorname{Tr}(\mathrm{Cov}(\text{vec}({\bm{H}}_{{\bm{Z}}})))\right|=\mathcal{O}(d^{2})$
@@
$\|\text{vec}({\bm{H}}_{{\bm{Z}}})\|=\mathcal{O}(d)$
holds with high probability. We now apply a result from [39, Theorem 5.44] concerning the spectral norm of a random matrix with independent heavy-tailed rows. This yields the stated bound on $\|\tilde{{\bm{H}}}\|$. Note that the assumption $n/d^{2}\in\mathbb{R}$ (from Assumption 4.1) ensures the validity of this application. ∎
@@
    $\|{\bm{w}}\|={\tilde{{\mathcal{O}}}}(1)$ and $\|{\bm{w}}\|_{\infty}={\tilde{{\mathcal{O}}}}(k^{-1/2})$,
@@
    $\|\tilde{{\bm{y}}}\|={\tilde{{\mathcal{O}}}}(n^{1/2})$ and $\|\tilde{{\bm{y}}}\|_{\infty}={\tilde{{\mathcal{O}}}}(1)$,
@@
(i) Follows from standard sub-Gaussian norm bounds, specifically [40, Proposition 2.5.2 and Theorem 3.1.1], in conjunction with Assumption 4.6.
(ii) The spectral norm bound on $\bm{F}$} is a direct consequence of the concentration of the spectral norm for sub-Gaussian random matrices; see [40, Theorem 4.4.5].
(iii) The bounds for $\tilde{{\bm{y}}}$ follow from the Gaussian concentration of Lipschitz functions [40, Theorem 5.2.2], noting that each element of $\tilde{{\bm{y}}}$ is defined via a Lipschitz transformation of sub-Gaussian variables. ∎
@@
    $\|\sigma_{\perp}^{\prime}({\bm{F}}\tilde{{\bm{H}}}^{T})\|\,/\,\|\tilde{{\bm{H}}}\|={\tilde{{\mathcal{O}}}}(k^{1/2})$,
@@
    $\|\sigma^{\prime}({\bm{F}}\tilde{{\bm{H}}}^{T})\|={\tilde{{\mathcal{O}}}}(k)$,
@@
    $\|{\bm{w}}^{T}\sigma({\bm{F}}\tilde{{\bm{H}}}^{T})\|_{\infty}\,/\,\|\tilde{{\bm{H}}}\|={\tilde{{\mathcal{O}}}}(1)$,
@@
(i) From Lemma 4.9, the elements of ${\bm{F}}\tilde{{\bm{H}}}^{T}$ converge in distribution to $\mathcal{N}(0,1)$ under the asymptotic regime specified by Assumption 4.1. Furthermore, the columns of ${\bm{F}}\tilde{{\bm{H}}}^{T}$ are independent by construction, although their covariance is not isotropic due to the structure imposed by $\mathrm{Cov}(\text{vec}({\bm{H}}_{{\bm{Z}}}))$. The function $\sigma_{\perp}^{\prime}(\cdot)$, being the centered component of the derivative, remains bounded and Lipschitz (by Assumption 4.8). Thus, the rows of $\sigma_{\perp}^{\prime}({\bm{F}}\tilde{{\bm{H}}}^{T})^{T}$ are independent anisotropic sub-Gaussian vectors. Using Gaussian concentration of Lipschitz functions [40, Theorem 5.2.2] and spectral norm bounds for matrices with independent sub-Gaussian rows [39, Theorem 5.39 and Eq. (5.26)], we obtain the desired result.
@@
$\|\sigma^{\prime}({\bm{F}}\tilde{{\bm{H}}}^{T})\|\leq\|\sigma_{\perp}^{\prime}({\bm{F}}\tilde{{\bm{H}}}^{T})\|+\alpha\|\bm{1}_{k\times d(d+1)}\|,$
where $\alpha:=\mathbb{E}_{z\sim\mathcal{N}(0,1)}[\sigma^{\prime}(z)]$ and $\bm{1}_{k\times d(d+1)}$ denotes a matrix of all ones. Since $\|\bm{1}_{k\times d(d+1)}\|=\mathcal{O}(k)$, the result follows from part (i).
(iii) The term ${\bm{w}}^{T}\sigma({\bm{F}}\tilde{{\bm{H}}}^{T})$ is a Lipschitz function of Gaussian random variables, and thus obeys Gaussian concentration [40, Theorem 5.2.2]. Since $\|{\bm{w}}\|={\tilde{{\mathcal{O}}}}(1)$ by Lemma C.2, the bound follows. ∎
@@
$\|\bm{\Delta}\|$ $=\left\|\frac{1}{n\sqrt{k}}\left({\bm{w}}\tilde{{\bm{y}}}^{T}\odot\sigma_{\perp}^{\prime}({\bm{F}}\tilde{{\bm{H}}}^{T})\right)\tilde{{\bm{H}}}-\frac{1}{nk}\left({\bm{w}}{\bm{w}}^{T}\sigma({\bm{F}}\tilde{{\bm{H}}}^{T})\odot\sigma^{\prime}({\bm{F}}\tilde{{\bm{H}}}^{T})\right)\tilde{{\bm{H}}}\right\|$, (S24)
$\leq\frac{1}{n\sqrt{k}}\left\|{\bm{w}}\tilde{{\bm{y}}}^{T}\odot\sigma_{\perp}^{\prime}({\bm{F}}\tilde{{\bm{H}}}^{T})\right\|\|\tilde{{\bm{H}}}\|+\frac{1}{nk}\left\|{\bm{w}}{\bm{w}}^{T}\sigma({\bm{F}}\tilde{{\bm{H}}}^{T})\odot\sigma^{\prime}({\bm{F}}\tilde{{\bm{H}}}^{T})\right\|\|\tilde{{\bm{H}}}\|$, (S25)
$\leq\frac{1}{n\sqrt{k}}\|{\bm{w}}\|_{\infty}\|\tilde{{\bm{y}}}\|_{\infty}\|\sigma_{\perp}^{\prime}({\bm{F}}\tilde{{\bm{H}}}^{T})\|\|\tilde{{\bm{H}}}\|+\frac{1}{nk}\|{\bm{w}}\|_{\infty}\|{\bm{w}}^{T}\sigma({\bm{F}}\tilde{{\bm{H}}}^{T})\|_{\infty}\|\sigma^{\prime}({\bm{F}}\tilde{{\bm{H}}}^{T})\|\|\tilde{{\bm{H}}}\|$, (S26)
@@
which holds with high probability, where $\beta\in[0,1]$ satisfies $\|\mathrm{Cov}(\text{vec}({\bm{H}}_{{\bm{Z}}}))\|={\tilde{{\mathcal{O}}}}(k^{-\beta+1})$ under Assumption 4.5.
@@
$\|{\bm{u}}{\bm{v}}^{T}\|$ $=\|{\bm{u}}\|\|{\bm{v}}\|=\frac{\alpha}{n\sqrt{k}}\|{\bm{w}}\|\|\tilde{{\bm{H}}}\|\|\tilde{{\bm{y}}}\|={\tilde{{\mathcal{O}}}}(k^{-\beta/2}),$ (S28)

### v2

@@ Appendix B Asymptotic distribution of ${\bm{F}}\text{vec}({\bm{H}}_{\bm{Z}})$
$\delta_{x}:=\frac{\left|\|{\bm{x}}_{\ell+1}\|_{2}^{2}-\operatorname{Tr}(\mathrm{Cov}({\bm{x}}_{\ell+1}))\right|}{\operatorname{Tr}(\mathrm{Cov}({\bm{x}}_{\ell+1}))},\quad\text{and}\quad\delta_{v}:=\frac{\left|\|{\bm{b}}\|_{2}^{2}-\operatorname{Tr}(\mathbb{E}[{\bm{b}}{\bm{b}}^{T}])\right|}{\operatorname{Tr}(\mathbb{E}[{\bm{b}}{\bm{b}}^{T}])}.$ (S16)
@@
$\frac{\left|\|\text{vec}({\bm{H}}_{\bm{Z}})\|_{2}^{2}-\operatorname{Tr}(\mathrm{Cov}(\text{vec}({\bm{H}}_{\bm{Z}})))\right|}{\operatorname{Tr}(\mathrm{Cov}(\text{vec}({\bm{H}}_{\bm{Z}})))}$ $=\frac{\left|\|{\bm{x}}_{\ell+1}\|_{2}^{2}\|{\bm{b}}\|_{2}^{2}-\operatorname{Tr}(\mathrm{Cov}({\bm{x}}_{\ell+1}))\operatorname{Tr}(\mathbb{E}[{\bm{b}}{\bm{b}}^{T}])\right|}{\operatorname{Tr}(\mathrm{Cov}({\bm{x}}_{\ell+1}))\operatorname{Tr}(\mathbb{E}[{\bm{b}}{\bm{b}}^{T}])}$ (S17)
@@
$t=\operatorname{Tr}(\mathrm{Cov}(\text{vec}({\bm{H}}_{{\bm{Z}}})\mid s=i))=\operatorname{Tr}(\mathrm{Cov}(\text{vec}({\bm{H}}_{{\bm{Z}}})\mid s=j))$
for any data source indices $i,j$. Furthermore, Corollary B.2 showed that $\|\text{vec}({\bm{H}}_{\bm{Z}})\|_{2}^{2}/t$ concentrates around 1 for all $s$ and $\bm{\xi}|s$. This leads us to the following unconditioned concentration result.
@@
$\left|\frac{\|\text{vec}({\bm{H}}_{{\bm{Z}}})\|_{2}^{2}}{t}-1\right|\leq\max_{s,\bm{\xi}|s}\left|\frac{\|\text{vec}({\bm{H}}_{{\bm{Z}}})\mid(s,\bm{\xi}|s)\|_{2}^{2}}{t}-1\right|=o(1),$ (S20)
@@
We apply a union bound over the (finite) set of data sources and integrate over the distribution of $\bm{\xi}|s$. Since the conditional deviation is uniformly bounded by $o(1)$, the same bound extends to the unconditioned case. ∎
@@
This completes our analysis of the concentration of $\|\text{vec}({\bm{H}}_{\bm{Z}})\|_{2}^{2}/t$ around 1, and thereby concludes the proof of Lemma 4.9.
@@ Appendix C Decomposition of the gradient matrix
where $\alpha:=\mathbb{E}_{z\sim\mathcal{N}(0,1)}[\sigma^{\prime}(z)]$ is the average slope, and $\sigma^{\prime}_{\perp}(z):=\sigma^{\prime}(z)-\alpha$ captures the zero-mean deviation. The decomposition is justified by the fact that the random variable entering $\sigma$—namely, ${\bm{f}}_{i}^{T}\tilde{{\bm{h}}}_{j}$—converges in distribution to $\mathcal{N}(0,1)$ for all $i,j$, by Lemma 4.9.
@@
This decomposition separates the gradient into two components: A spike term ${\bm{u}}{\bm{v}}^{T}$, which is expected to dominate, and a residual term $\bm{\Delta}$.
Our next goal is to establish that $\|{\bm{u}}{\bm{v}}^{T}\|_{2}\gg\|\bm{\Delta}\|_{2}$, thereby confirming that the spike term governs the spectral structure of $\bm{G}$}.
@@
$\|\tilde{{\bm{H}}}\|_{2}\,/\,\|\mathrm{Cov}(\text{vec}({\bm{H}}_{{\bm{Z}}}))\|_{2}^{1/2}={\tilde{{\mathcal{O}}}}(d).$ (S23)
@@
$\left|\|\text{vec}({\bm{H}}_{{\bm{Z}}})\|_{2}^{2}-\operatorname{Tr}(\mathrm{Cov}(\text{vec}({\bm{H}}_{{\bm{Z}}})))\right|=\mathcal{O}(d^{2})$
@@
$\|\text{vec}({\bm{H}}_{{\bm{Z}}})\|_{2}=\mathcal{O}(d)$
holds with high probability. We now apply a result from [38, Theorem 5.44] concerning the spectral norm of a random matrix with independent heavy-tailed rows. This yields the stated bound on $\|\tilde{{\bm{H}}}\|_{2}$. Note that the assumption $n/d^{2}\in\mathbb{R}$ (from Assumption 4.1) ensures the validity of this application. ∎
@@
    $\|{\bm{w}}\|_{2}={\tilde{{\mathcal{O}}}}(1)$ and $\|{\bm{w}}\|_{\infty}={\tilde{{\mathcal{O}}}}(k^{-1/2})$,
@@
    $\|\tilde{{\bm{y}}}\|_{2}={\tilde{{\mathcal{O}}}}(n^{1/2})$ and $\|\tilde{{\bm{y}}}\|_{\infty}={\tilde{{\mathcal{O}}}}(1)$,
@@
(i) Follows from standard sub-Gaussian norm bounds, specifically [39, Proposition 2.5.2 and Theorem 3.1.1], in conjunction with Assumption 4.6.
(ii) The spectral norm bound on $\bm{F}$} under Assumption 4.6 is a direct consequence of the concentration of the spectral norm for sub-Gaussian random matrices; see [39, Theorem 4.4.5].
(iii) The bounds for $\tilde{{\bm{y}}}$ follow from the Gaussian concentration of Lipschitz functions [39, Theorem 5.2.2], noting that each element of $\tilde{{\bm{y}}}$ is defined via a Lipschitz transformation of sub-Gaussian variables. ∎
@@
    $\|\sigma_{\perp}^{\prime}({\bm{F}}\tilde{{\bm{H}}}^{T})\|_{2}\,/\,\|\tilde{{\bm{H}}}\|_{2}={\tilde{{\mathcal{O}}}}(1)$,
@@
    $\|\sigma^{\prime}({\bm{F}}\tilde{{\bm{H}}}^{T})\|_{2}={\tilde{{\mathcal{O}}}}(k)$,
@@
    $\|{\bm{w}}^{T}\sigma({\bm{F}}\tilde{{\bm{H}}}^{T})\|_{\infty}\,/\,\|\tilde{{\bm{H}}}\|_{2}={\tilde{{\mathcal{O}}}}(k^{-1/2})$,
@@
(i) Condition on $\tilde{{\bm{H}}}$. By construction, the rows of ${\bm{F}}\tilde{{\bm{H}}}^{T}$ are independent centered Gaussian vectors (with generally anisotropic covariance inherited from $\mathrm{Cov}(\text{vec}({\bm{H}}_{{\bm{Z}}}))$). The function $\sigma_{\perp}^{\prime}(\cdot)$ is bounded and Lipschitz by Assumption 4.8 so we apply Gaussian concentration for Lipschitz functions [39, Theorem 5.2.2]. This gives us the rows of $\sigma_{\perp}^{\prime}({\bm{F}}\tilde{{\bm{H}}}^{T})$ as independent sub-Gaussian vectors with sub-Gaussian norm proportional to $\|\tilde{{\bm{H}}}\|_{2}\operatorname{Tr}(\mathrm{Cov}(\text{vec}({\bm{H}}_{{\bm{Z}}})))^{-1/2}$ with $\operatorname{Tr}(\mathrm{Cov}(\text{vec}({\bm{H}}_{{\bm{Z}}})))\asymp k$ by Assumption 4.4. Finally, employing the spectral norm bound for matrices with independent sub-Gaussian rows [38, Theorem 5.39 and Eq. (5.26)], and then removing the conditioning, yields the stated bound.
@@
$\|\sigma^{\prime}({\bm{F}}\tilde{{\bm{H}}}^{T})\|_{2}\leq\|\sigma_{\perp}^{\prime}({\bm{F}}\tilde{{\bm{H}}}^{T})\|_{2}+\alpha\|\bm{1}_{k\times n}\|_{2},$
where $\alpha:=\mathbb{E}_{z\sim\mathcal{N}(0,1)}[\sigma^{\prime}(z)]$ and $\bm{1}_{k\times n}$ denotes a matrix of all ones. Since $\|\bm{1}_{k\times n}\|_{2}=\sqrt{kn}={\tilde{{\mathcal{O}}}}(k)$ under Assumption 4.1, the result follows from part (i).
(iii) Similar to (i), using Gaussian concentration of Lipschitz functions [39, Theorem 5.2.2] along with $\|{\bm{w}}\|_{2}={\tilde{{\mathcal{O}}}}(1)$ from Lemma C.2 gives the claim. ∎
@@
$\|\bm{\Delta}\|_{2}$ $=\left\|\frac{1}{n\sqrt{k}}\left({\bm{w}}\tilde{{\bm{y}}}^{T}\odot\sigma_{\perp}^{\prime}({\bm{F}}\tilde{{\bm{H}}}^{T})\right)\tilde{{\bm{H}}}-\frac{1}{nk}\left({\bm{w}}{\bm{w}}^{T}\sigma({\bm{F}}\tilde{{\bm{H}}}^{T})\odot\sigma^{\prime}({\bm{F}}\tilde{{\bm{H}}}^{T})\right)\tilde{{\bm{H}}}\right\|_{2},$ (S24)
$\leq\frac{1}{n\sqrt{k}}\left\|{\bm{w}}\tilde{{\bm{y}}}^{T}\odot\sigma_{\perp}^{\prime}({\bm{F}}\tilde{{\bm{H}}}^{T})\right\|_{2}\|\tilde{{\bm{H}}}\|_{2}+\frac{1}{nk}\left\|{\bm{w}}{\bm{w}}^{T}\sigma({\bm{F}}\tilde{{\bm{H}}}^{T})\odot\sigma^{\prime}({\bm{F}}\tilde{{\bm{H}}}^{T})\right\|_{2}\|\tilde{{\bm{H}}}\|_{2},$ (S25)
$\leq\frac{1}{n\sqrt{k}}\|{\bm{w}}\|_{\infty}\|\tilde{{\bm{y}}}\|_{\infty}\|\sigma_{\perp}^{\prime}({\bm{F}}\tilde{{\bm{H}}}^{T})\|_{2}\|\tilde{{\bm{H}}}\|_{2}+\frac{1}{nk}\|{\bm{w}}\|_{\infty}\|{\bm{w}}^{T}\sigma({\bm{F}}\tilde{{\bm{H}}}^{T})\|_{\infty}\|\sigma^{\prime}({\bm{F}}\tilde{{\bm{H}}}^{T})\|_{2}\|\tilde{{\bm{H}}}\|_{2},$ (S26)
@@
which holds with high probability, where $\beta\in(0,1)$ satisfies $\|\mathrm{Cov}(\text{vec}({\bm{H}}_{{\bm{Z}}}))\|_{2}={\mathcal{O}}(k^{-\beta+1})$ and $\eta=o(k^{\beta})$ under Assumption 4.5.
@@
$\|{\bm{u}}{\bm{v}}^{T}\|_{2}$ $=\|{\bm{u}}\|_{2}\|{\bm{v}}\|_{2}=\frac{\alpha}{n\sqrt{k}}\|{\bm{w}}\|_{2}\|\tilde{{\bm{H}}}\|_{2}\|\tilde{{\bm{y}}}\|_{2}={\tilde{{\mathcal{O}}}}(k^{-\beta/2}),$ (S28)